# Time-asymptotic self-similarity of the damped compressible Euler equations in parabolic scaling variables

### v1

@@ 1 Introduction
Consider the one-dimensional flow of an isentropic gas subject to external friction, such as a flow through a porous medium, described by the damped Euler equations. We first consider these equations on the real line, given by
$\widetilde{\rho}_{t}+\widetilde{m}_{x}$ $=0,$ (1.1)
$\widetilde{m}_{t}+\Big{(}\frac{\widetilde{m}^{2}}{\widetilde{\rho}}+\mathfrak{p}(\widetilde{\rho})\Big{)}_{x}+\alpha\widetilde{m}$ $=0,$
@@
$\widetilde{\rho}_{t}+\widetilde{m}_{x}$ $=0,$
$\mathfrak{p}(\widetilde{\rho})_{x}+\alpha\widetilde{m}$ $=0,$
@@
In the literature, such self-similar solutions $(\widetilde{\rho}_{\mathrm{se}},\widetilde{m}_{\mathrm{se}})$ are usually called nonlinear diffusion waves, as for example in [CY18]. In the work [HsL93], a uniform convergence rate was shown under increased regularity assumptions on the solution $(\rho,m)$. In [HsL92], the same authors derived similar results for the damped $p$-system, which is a formulation of system (1.1) in Lagrangian coordinates. Since then, these results have been extended in various ways, see [Nis96, SeX97, NWY00, HuP03, Mei09, GeT19] and references therein, especially by derivation of improved convergence rates but also by allowing for more general solution classes. The equations have also been studied subject to a time-dependent damping, see for instance [CY18, GHW22, GH23]. In the case of finite mass, we also refer to [HMP05, HPW11, GeH19], where the convergence towards the famous Barenblatt solution is shown.
@@
This approach is in accordance with [MiS24], where the long-time behavior of solutions to a reaction-diffusion system was studied. The scaling (1.7) emphasizes the time-asymptotic parabolic behavior of the system, and it was already used by Gallay and Raugel [GaR98, GaR97, GaR00] to analyze the asymptotic behavior of scalar damped hyperbolic equations. Nevertheless, to the authors’ knowledge, the present paper is the first one to adapt this method to a damped hyperbolic system and to combine it with a relative entropy approach.
@@
where the term $\frac{1}{2}{\mathcal{E}}(\tau)$ arises from the drift terms $\frac{y}{2}\rho_{y}$ and $\frac{y}{2}n_{y}$ in (1.8) and is thus due to the parabolic scaling. Moreover, ${\mathcal{D}}_{\alpha}\geq 0$ is a dissipative term associated with the frictional damping, while the term $\Xi$ does not have a clear sign and has to be controlled by the other terms. To this end, the term $\frac{1}{2}{\mathcal{E}}(\tau)$ is necessary, and its appearance can be considered the main effect of the transformation into scaling variables. Moreover, it finally leads to the exponential decay of ${\mathcal{E}}(\tau)$ as $\tau\to\infty$ by means of Gronwall’s inequality. Since the limiting profile is not a true steady state if we consider different limits at spatial infinity, the function ${\mathcal{E}}$ cannot be a true Lyapunov function. Instead, we only show that ${\mathcal{E}}$ becomes a Lyapunov function for large $\tau$, which is sufficient for the convergence and comparable with the results in [MiS24].
Notice that smooth solutions to (1.8) even satisfy (1.12) as an equality. Since those exist only up to a finite time in general, we consider a framework of weak solutions instead, where results on the global existence of solutions satisfying an entropy inequality are available in the one-dimensional case, see [DCL89, HuP06]. However, in the multi-dimensional case, analogous results seem to be out of reach for weak solutions, which is why we instead study a further relaxed solvability concept in this setting, namely energy-variational solutions. This concept was recently introduced for a general class of hyperbolic conservation laws in [EiL24], including the compressible Euler equations, and has further been investigated for other models, see [ALR24, EHL23, LaR23, Las24, Las25]. We establish the global existence of energy-variational solutions and show that our results on the long-time behavior of weak solutions can be transferred to this case.
@@ 2 Convergence results in one dimension
Notice that one can also derive this weak formulation by first taking the weak formulations of the unscaled equations (1.1) and then transforming them into parabolic coordinates.
@@
Inequality (2.8) is a weak formulation of (2.3) relaxed to an inequality. It can equivalently be derived by transforming an analogously relaxed version of (1.5) into parabolic scaling variables $(\tau,y)$. The local entropy inequality (2.8) is a classical assumption, and weak solutions satisfying (2.8) are usually called weak entropy solutions. Moreover, the additional boundedness assumptions on $\rho$ and $\frac{n}{\rho}$ in (E1) seem to be necessary for the derivation of an inequality for the total relative entropy. Observe that the global existence of weak solutions satisfying (E1) was shown in [DCL89] for $\mathfrak{p}(\rho)=k\rho^{\gamma}$ with $21<\gamma<3/$ 2, which was extended in [HuP06] to the case $\gamma<3$.
@@
$\alpha\int_{\tau}^{\infty}\!\!\!\int_{\mathbb{R}}\rho\,\Big{|}\frac{n}{\rho}-\frac{n^{*}}{\rho^{*}}\Big{|}^{2}\,\mathrm{d}y\,\mathrm{d}\tau^{\prime}\leq\mathrm{e}^{-(\frac{1}{2}{-}\theta)\tau+\frac{1}{2}\mu}\Big{(}{\mathcal{E}}(0;\rho_{0},n_{0}\,|\,\rho^{*},n^{*})+\frac{K}{\theta}\Big{)}+\frac{K}{2}\mathrm{e}^{-\tau/2}$ (2.13)
@@
Let $\alpha\geq 0$ and $\rho_{+}=\rho_{-}\equiv\hat{\rho}>0$, and let $\mathfrak{p}\in C([0,\infty))\cap C^{1}(0,\infty)$ with $\mathfrak{p}^{\prime}>0$ in $(0,\infty)$. Consider initial data $(\rho_{0},n_{0})\in L^{1}_{\mathrm{loc}}({\mathbb{R}})^{2}$ such that $\rho_{0}\geq 0$ and ${\mathcal{E}}(0;\rho_{0},n_{0}\,|\,\hat{\rho},0)<\infty$. Let $(\rho,n)$ be a weak solution to (1.8) satisfying (E1) or (E2). Then
@@
$\alpha\int_{t}^{\infty}\!\!\!\int_{\mathbb{R}}\widetilde{\rho}(t,y\sqrt{1{+}t^{\prime}})\,\bigg{|}\frac{\widetilde{m}(t^{\prime},y\sqrt{1{+}t^{\prime}})}{\widetilde{\rho}(t^{\prime},y\sqrt{1{+}t^{\prime}})}-\frac{\widetilde{m}_{\mathrm{se}}(t^{\prime},y\sqrt{1{+}t^{\prime}})}{\widetilde{\rho}_{\mathrm{se}}(t^{\prime},y\sqrt{1{+}t^{\prime}})}\bigg{|}^{2}\,\mathrm{d}x\,\mathrm{d}t^{\prime}\leq C(1{+}t)^{-(\frac{1}{2}-\theta)},$
which provides an additional rate for the convergence of time-averaged kinetic energy as $t\to\infty$.
@@ 3.1 Preparatory results
Let $\mathfrak{p}\in C([0,\infty))\cap C^{1}(0,\infty)$ with $\mathfrak{p}^{\prime}>0$ in $(0,\infty)$. Let $0<\delta<M$. Then there exists $C>0$ such that
@@
First consider $(\rho_{1},n_{1})=(0,0)$ and $(\rho_{2},n_{2})\in\{(0,0)\}\cup[\delta,M]\times{\mathbb{R}}$. Then we have
@@
To estimate the latter term, consider first the case $20<\rho_{1}<\delta/$ 2. We use (1.3) to deduce $\rho_{1}\big{(}h^{\prime}(\rho_{1})-h^{\prime}(\rho_{2})\big{)}=h(\rho_{1}\,|\,\rho_{2})+\mathfrak{p}(\rho_{1})-\mathfrak{p}(\rho_{2})$. First, we have $\big{|}\mathfrak{p}(\rho_{1})-\mathfrak{p}(\rho_{2})\big{|}\leq\mathfrak{p}(M)-\mathfrak{p}(\delta/2)$ by monotonicity of $\mathfrak{p}$. Second, we use Taylor’s theorem and $h^{\prime\prime}(z)>0$ for $z>0$ to conclude
@@
$\rho-\rho_{1}=(\sqrt{\rho}-\sqrt{\rho_{1}})^{2}+2\sqrt{\rho_{1}}(\sqrt{\rho}-\sqrt{\rho_{1}})\in L^{1}(Q)+L^{2}(Q)$
@@
yields $\mathrm{e}^{-\tau/2}(n-n_{1})\mathrm{e}^{-\tau/2}(n-n_{1})\in L^{1}(Q)+L^{2}(Q)$. In view of the assumptions, we conclude the integrability of the right-hand side of (3.10) in both cases, which completes the proof. ∎
@@
$\widetilde{\eta}(\tau;\rho,n\,|\,\overline{\rho},0)\geq\begin{cases}C\big{(}\mathrm{e}^{-\tau}|n|^{2}+|\rho-\overline{\rho}|^{2}\big{)}&\text{ if }\rho\in[0,r_{0}],\\ C\big{(}(\mathrm{e}^{-\tau}|n|^{2})^{\gamma/(\gamma+1)}+|\rho-\overline{\rho}|^{\gamma}\big{)}&\text{ if }\rho\in(r_{0},\infty),\end{cases}$ (3.11)
@@
for all $\overline{\rho}\in K$. Indeed, by Taylor’s formula, we have for $\rho\in K$ that
@@
$\widetilde{\eta}(\rho,n\,|\,\overline{\rho},0)\geq\mathrm{e}^{-\tau}\frac{|n|^{2}}{2r_{0}}+c|\rho-\overline{\rho}|^{2}\qquad\text{if }\rho\leq r_{0}.$
@@
$\widetilde{\eta}(\rho,n\,|\,\overline{\rho},0)\geq C_{\gamma}(\mathrm{e}^{-\tau}|n|^{2})^{\gamma/(\gamma+1)}\Big{(}\frac{1}{2}\Big{)}^{\gamma/(\gamma+1)}+\frac{c}{2}|\rho-\overline{\rho}|^{\gamma}\qquad\text{if }\rho>r_{0}.$
@@ 3.2 Weak solutions satisfying (E1)
Let $\mathfrak{p}\in C([0,\infty))\cap C^{1}(0,\infty)$ with $\mathfrak{p}^{\prime}>0$, and let $(\rho,n)$ be a weak solution to (1.8) fulfilling (E1) with initial data $(\rho_{0},n_{0})\in L^{1}_{\mathrm{loc}}({\mathbb{R}})^{2}$. Let $(\overline{\rho},\overline{n})\in C^{1}((0,\infty)\times{\mathbb{R}})^{2}\cap C([0,\infty)\times{\mathbb{R}})^{2}$ be bounded such that $\inf\overline{\rho}>0$, or let $(\overline{\rho},\overline{n})\equiv(0,0)$ and additionally assume that $\mathfrak{p}$ satisfies (2.14). Let
@@
As observed in Remark 3.10, criterion (E2) is satisfied, so that ${\mathcal{E}}(\tau;\rho,n\,|\,\rho_{\infty},0)\in L_{\mathrm{loc}}^{\infty}([0,\infty))$ for some $\rho_{\infty}\in C^{1}({\mathbb{R}})$ satisfying (2.7). From (3.21) and Proposition 3.3 we further obtain ${\mathcal{E}}(\tau;\rho,n\,|\,\overline{\rho},\overline{n})\in L_{\mathrm{loc}}^{\infty}([0,\infty))$. Using now (3.21) and the estimates from Lemmas 3.5, 3.6 and 3.7, we obtain $\xi\in L^{1}_{\mathrm{loc}}([0,\infty);L^{1}({\mathbb{R}}))$. From the finiteness of ${\mathcal{E}}(0;\rho_{0},n_{0}\,|\,\hat{\rho},0)$ and ${\mathcal{E}}(0;\rho_{0},n_{0}\,|\,\overline{\rho}(0),\overline{n}(0))$, we further conclude ${\mathcal{E}}(0;\rho_{0},n_{0}\,|\,\overline{\rho}(0),\overline{n}(0))<\infty$ by Proposition 3.3. Now the statement follows from Theorem 3.9. ∎

### v2

@@ 1 Introduction
Consider the flow of an isentropic gas subject to external friction, such as a flow through a porous medium, described by the damped Euler equations. We first consider these equations on the real line, given by
$\widetilde{\rho}_{t}+\widetilde{m}_{x}$ =0=0,= 0 , (1.1)
$\widetilde{m}_{t}+\Big{(}\frac{\widetilde{m}^{2}}{\widetilde{\rho}}+\mathfrak{p}(\widetilde{\rho})\Big{)}_{x}+\alpha\widetilde{m}$ =0=0,= 0 ,
@@
$\widetilde{\rho}_{t}+\widetilde{m}_{x}$ =0=0,= 0 ,
$\mathfrak{p}(\widetilde{\rho})_{x}+\alpha\widetilde{m}$ =0=0,= 0 ,
@@
In the literature, such self-similar solutions $(\widetilde{\rho}_{\mathrm{se}},\widetilde{m}_{\mathrm{se}})$ are usually called nonlinear diffusion waves, as for example in [CY18]. In the work [HsL93], a uniform convergence rate was shown under increased regularity assumptions on the solution $(\rho,m)$. In [HsL92], the same authors derived similar results for the damped $p$-system, which is a formulation of system (1.1) in Lagrangian coordinates. Since then, these results have been extended in various ways, see [Nis96, SeX97, NWY00, HuP03, Mei09, GeT19] and references therein, especially by derivation of improved convergence rates but also by allowing for more general solution classes. The equations have also been studied subject to a time-dependent damping, see for instance [CY18, GHW22, GH23]. In the case of finite mass, we refer to [HMP05, HPW11, GeH19], where the convergence towards the famous Barenblatt solution is shown.
@@
This approach is in accordance with [MiS24], where the long-time behavior of solutions to a reaction-diffusion system was studied. The scaling (1.7) emphasizes the time-asymptotic parabolic behavior of the system, and it was already used by Gallay and Raugel [GaR98, GaR97, GaR00] to analyze the asymptotic behavior of scalar damped hyperbolic equations. To the authors’ knowledge, the present paper is the first one to adapt this method to a damped hyperbolic system and to combine it with a relative entropy approach.
@@
where the term $\frac{1}{2}{\mathcal{E}}(\tau)$ arises from the drift terms $\frac{y}{2}\rho_{y}$ and $\frac{y}{2}n_{y}$ in (1.8) and is thus due to the parabolic scaling. Moreover, ${\mathcal{D}}_{\alpha}\geq 0$ is a dissipative term associated with the frictional damping, while the term $\Xi$ does not have a clear sign and has to be controlled by the other terms. To this end, the term $\frac{1}{2}{\mathcal{E}}(\tau)$ is necessary, and its appearance can be considered the main effect of the transformation into scaling variables. Moreover, it finally leads to the exponential decay of ${\mathcal{E}}(\tau)$ as $\tau\to\infty$ by means of Gronwall’s inequality. However, since the limiting profile is not a true steady state if we consider different limits at spatial infinity, the function ${\mathcal{E}}$ cannot be a true Lyapunov function. Instead, we only show that ${\mathcal{E}}$ becomes a Lyapunov function for large $\tau$, which is sufficient for the convergence and comparable with the results in [MiS24].
Notice that smooth solutions to (1.8) even satisfy (1.12) as an equality. Since those exist only up to a finite time in general, we consider a framework of weak solutions instead, where results on the global existence of solutions satisfying an entropy inequality are available in the one-dimensional case, see [DCL89, HuP06]. However, in the multi-dimensional case, analogous results for weak solutions seem to be out of reach, which is why we instead study a further relaxed solvability concept in this setting, namely energy-variational solutions. This concept was recently introduced for a general class of hyperbolic conservation laws in [EiL24], including the compressible Euler equations, and has further been investigated for other models, see [ALR24, EHL23, LaR23, Las24, Las25]. We establish the global existence of energy-variational solutions and show that our results on the long-time behavior of weak solutions can be transferred to this case.
@@ 2 Convergence results in one dimension
Notice that one can also derive this weak formulation by transforming the weak formulations of the unscaled equations (1.1) into parabolic coordinates.
@@
Inequality (2.8) is a weak formulation of (2.3) relaxed to an inequality. It can equivalently be derived by transforming an analogously relaxed version of (1.5) into parabolic scaling variables $(\tau,y)$. The local entropy inequality (2.8) is a classical assumption, and weak solutions satisfying (2.8) are usually called weak entropy solutions. Moreover, the additional boundedness assumptions on $\rho$ and $\frac{n}{\rho}$ in (E1) seem to be necessary for the derivation of an inequality for the total relative entropy. Observe that the global existence of weak solutions satisfying (E1) was shown in [DCL89] for $\mathfrak{p}(\rho)=k\rho^{\gamma}$ with $1<\gamma<3/2$, which was extended in [HuP06] to the case $\gamma<3$.
@@
$\alpha\int_{\tau}^{\infty}\!\!\!\int_{\mathbb{R}}\rho\,\Big{|}\frac{n}{\rho}-\frac{n^{*}}{\rho^{*}}\Big{|}^{2}\,\mathrm{d}y\,\mathrm{d}\tau^{\prime}\leq\mathrm{e}^{-(\frac{1}{2}{-}\theta)\tau+\frac{1}{2}\mu}\Big{(}{\mathcal{E}}(0;\rho_{0},n_{0}\,|\,\rho^{*},n^{*})+\frac{K}{\theta}\Big{)}+2K\mathrm{e}^{-\tau/2}$ (2.13)
@@
Let $\alpha\geq 0$ and $\rho_{+}=\rho_{-}\equiv\hat{\rho}>0$, and let $\mathfrak{p}\in C([0,\infty))\cap C^{1}(0,\infty)$ with $\mathfrak{p}^{\prime}>0$ in (0,)(0$\infty$)( 0 , ∞ ). Consider initial data $(\rho_{0},n_{0})\in L^{1}_{\mathrm{loc}}({\mathbb{R}})^{2}$ such that $\rho_{0}\geq 0$ and ${\mathcal{E}}(0;\rho_{0},n_{0}\,|\,\hat{\rho},0)<\infty$. Let $(\rho,n)$ be a weak solution to (1.8) satisfying (E1) or (E2). Then
@@
$\alpha\int_{t}^{\infty}\!\!\!\int_{\mathbb{R}}\widetilde{\rho}(t,y\sqrt{1{+}t^{\prime}})\,\bigg{|}\frac{\widetilde{m}(t^{\prime},y\sqrt{1{+}t^{\prime}})}{\widetilde{\rho}(t^{\prime},y\sqrt{1{+}t^{\prime}})}-\frac{\widetilde{m}_{\mathrm{se}}(t^{\prime},y\sqrt{1{+}t^{\prime}})}{\widetilde{\rho}_{\mathrm{se}}(t^{\prime},y\sqrt{1{+}t^{\prime}})}\bigg{|}^{2}\,\mathrm{d}y\,\mathrm{d}t^{\prime}\leq C(1{+}t)^{-(\frac{1}{2}-\theta)},$
which provides an additional rate for the convergence of the time-averaged kinetic energy as $t\to\infty$.
@@ 3.1 Preparatory results
Let $\mathfrak{p}\in C([0,\infty))\cap C^{1}(0,\infty)$ with $\mathfrak{p}^{\prime}>0$ in (0,)(0$\infty$)( 0 , ∞ ). Let $0<\delta<M$. Then there exists $C>0$ such that
@@
First consider $(\rho_{1},n_{1})=(0,0)$ and $(\rho_{2},n_{2})\in[\delta,M]\times{\mathbb{R}}$. Then we have
@@
To estimate the latter term, consider first the case $0<\rho_{1}<\delta/2$. We use (1.3) to deduce $\rho_{1}\big{(}h^{\prime}(\rho_{1})-h^{\prime}(\rho_{2})\big{)}=h(\rho_{1}\,|\,\rho_{2})+\mathfrak{p}(\rho_{1})-\mathfrak{p}(\rho_{2})$. First, we have $\big{|}\mathfrak{p}(\rho_{1})-\mathfrak{p}(\rho_{2})\big{|}\leq\mathfrak{p}(M)-\mathfrak{p}(\delta/2)$ by monotonicity of $\mathfrak{p}$. Second, we use Taylor’s theorem and $h^{\prime\prime}(z)>0$ for $z>0$ to conclude
@@
$\rho-\rho_{1}=(\sqrt{\rho}-\sqrt{\rho_{1}})^{2}+2\sqrt{\rho_{1}}(\sqrt{\rho}-\sqrt{\rho_{1}})\in L^{1}(Q)+L^{2}(Q),$
@@
yields $\mathrm{e}^{-\tau/2}(n-n_{1})\in L^{1}(Q)+L^{2}(Q)$. In view of the assumptions, we conclude the integrability of the right-hand side of (3.10) in both cases, which completes the proof. ∎
@@
$\eta(\tau;\rho,n\,|\,\overline{\rho},0)\geq\begin{cases}C\big{(}\mathrm{e}^{-\tau}|n|^{2}+|\rho-\overline{\rho}|^{2}\big{)}&\text{ if }\rho\in[0,r_{0}],\\ C\big{(}(\mathrm{e}^{-\tau}|n|^{2})^{\gamma/(\gamma+1)}+|\rho-\overline{\rho}|^{\gamma}\big{)}&\text{ if }\rho\in(r_{0},\infty),\end{cases}$ (3.11)
@@
for all $\overline{\rho}\in K$. To this end, we show that
$(\rho,\overline{\rho})\mapsto\frac{|\rho-\overline{\rho}|^{2}}{h(\rho\,|\,\overline{\rho})}\chi_{[0,r_{0}]}(\rho)+\frac{h(\rho-\overline{\rho})}{h(\rho\,|\,\overline{\rho})}\chi_{(r_{0},\infty)}(\rho)$
defines a bounded function on $(0,\infty)\times K$, where $\chi_{M}$ is the characteristic function of a set $M\subset{\mathbb{R}}$. Indeed, by Taylor’s formula, we have for $\rho\in K$ that
@@
$\eta(\tau;\rho,n\,|\,\overline{\rho},0)\geq\mathrm{e}^{-\tau}\frac{|n|^{2}}{2r_{0}}+c|\rho-\overline{\rho}|^{2}\qquad\text{if }\rho\leq r_{0}.$
@@
$\eta(\tau;\rho,n\,|\,\overline{\rho},0)\geq C_{\gamma}(\mathrm{e}^{-\tau}|n|^{2})^{\gamma/(\gamma+1)}\Big{(}\frac{1}{2}\Big{)}^{\gamma/(\gamma+1)}+\frac{c}{2}|\rho-\overline{\rho}|^{\gamma}\qquad\text{if }\rho>r_{0}.$
@@ 3.2 Weak solutions satisfying (E1)
Let $\mathfrak{p}\in C([0,\infty))\cap C^{1}(0,\infty)$ with $\mathfrak{p}^{\prime}>0$, and let $(\rho,n)$ be a weak solution to (1.8) fulfilling (E1) with initial data $(\rho_{0},n_{0})\in L^{1}_{\mathrm{loc}}({\mathbb{R}})^{2}$. Let either $(\overline{\rho},\overline{n})\in C^{1}((0,\infty)\times{\mathbb{R}})^{2}\cap C([0,\infty)\times{\mathbb{R}})^{2}$ be bounded such that $\inf\overline{\rho}>0$, or let $(\overline{\rho},\overline{n})\equiv(0,0)$ and assume that $\mathfrak{p}$ satisfies (2.14). Let
@@
As observed in Remark 3.10, criterion (E2) is satisfied, so that ${\mathcal{E}}(\tau;\rho,n\,|\,\rho_{\infty},0)\in L_{\mathrm{loc}}^{\infty}([0,\infty))$ for some $\rho_{\infty}\in C^{1}({\mathbb{R}})$ satisfying (2.7). From (3.21) and Proposition 3.3 we further obtain ${\mathcal{E}}(\tau;\rho,n\,|\,\overline{\rho},\overline{n})\in L_{\mathrm{loc}}^{\infty}([0,\infty))$. Using now (3.21) and the estimates from Lemmas 3.5, 3.6 and 3.7, we obtain $\xi\in L^{1}_{\mathrm{loc}}([0,\infty);L^{1}({\mathbb{R}}))$. From the finiteness of ${\mathcal{E}}(0;\rho_{0},n_{0}\,|\,\hat{\rho},0)$ and ${\mathcal{E}}(0;\overline{\rho}(0),\overline{n}(0)\,|\,\hat{\rho},0)$, we further conclude ${\mathcal{E}}(0;\rho_{0},n_{0}\,|\,\overline{\rho}(0),\overline{n}(0))<\infty$ by Proposition 3.3. Now the statement follows from Theorem 3.9. ∎